OpenStudy (anonymous):

Find dy/dx in terms of y when x and y are related by the following equation: x=y−y^3

OpenStudy (anonymous):

This is an implicit differentiation problem with a twist :D But, more on that later... Could you please differentiate both sides of the equation with respect to x? (Pretty Please? :) )

OpenStudy (anonymous):

is this right?\[1=1-3*y^2\]

OpenStudy (anonymous):

Naaaahh :D You can differentiate y the way you differentiate x, but ALWAYS remember to place a dy/dx afterwards... remember, we're differentiating with respect to x, not y ;)

OpenStudy (anonymous):

Here, I'll give you an example to follow :) Say we have to differentiate \[1=x^2+y^2\]with respect to x. That's easy, we differentiate the left (which is just a constant anyway) and the right... \[0=2x+2y\frac{dy}{dx}\] See the difference? With x^2, we differentiate normally, but with y^2, we sorta differentiate normally, except when we're done, we put a dy/dx after it :) Now you try :>

OpenStudy (anonymous):

1/1-3y^2

OpenStudy (anonymous):

My connection sucks, but wouldn't it be 1=dy/dx-(3y²)dy/dx?

OpenStudy (anonymous):

and then taking dy/dx as a common factor: 1=dy/dx(1-3y²)

OpenStudy (anonymous):

Yeah, that's it, and then you can factor out that dy/dx, like this \[1=\frac{dy}{dx}(1-3y^2)\] And now, it's just a matter of isolating dy/dx, which...

OpenStudy (anonymous):

dy/dx=1/(1-3y²)

OpenStudy (anonymous):

as @AQEE had already mentioned :) But it's good that you managed to arrive at the answer yourself :D Nice job ^.^

OpenStudy (anonymous):

thanks @PeterPan !

OpenStudy (anonymous):

No problem :>